ronaldneedanswers:

Which statements are true for the given geometric sequence? Check all that apply. The domain is the set of natural numbers. The range is the set of natural numbers. The recursive formula representing the sequence is f(x + 1) = Three-halves(f(x )) when f(1) = 4. An explicit formula representing the sequence is f(x) = 4(three-halves) Superscript x The sequence shows exponential growth.

Vocaloid:

What’s the geometric sequence? (What numbers do they give you?)

KingJay22381:

The statements are true for the given geometric sequence. The domain is the set of natural numbers. The recursive formula representing the sequence is f(x + 1) = 3/2(f(x )) when f(1) = 4. The sequence shows exponential growth. We have given the graph. What is the geometric sequence? A geometric sequence is a sequence in which the ratio of every two successive terms is a constant. The graph is given below we can see that The domain is the set of natural numbers. The recursive formula representing the sequence is f(x + 1) = 3/2(f(x )) when f(1) = 4. The given graph grows exponentially. The sequence shows exponential growth